Computing Revenue and Gross Profit on Long-term Construction Contract

Supplier Corp. enters into a government contract in 2020 to provide computer equipment for $2 million. The contract consists of a single performance obligation to provide specified equipment in three years. Total costs estimated by Supplier Corp. for the contract are $1.4 million. The equipment is highly specialized and has no alternative uses. As negotiated in the contract, any costs incurred by Supplier Corp. plus a specified profit margin will be paid to Supplier Corp. in the event of a contract cancellation. Actual costs incurred in 2020 were $650,000 including unexpected cost overruns of $80,000 due to labor inefficiencies.

At the end of 2021, the estimate of total costs has increased to $1.50 million due to an increase in cost of materials. Actual costs incurred to date are $1,150,000, excluding year one inefficiencies.

a. Calculate (1) recognized revenue, (2) the gross profit, and (3) adjusted contract margin to be recorded in 2021.

Note: Round each of your answers to the nearest dollar.

1. Recognized revenue Answer
2. Gross profit Answer
3. Adjusted contract margin Answer

b. Calculate (1) cumulative recognized revenue, (2) cumulative gross profit, and (3) cumulative adjusted contract margin at the end of 2021.

Note: Round each of your answers to the nearest dollar.

1. Cumulative recognized revenue Answer
2. Cumulative gross profit Answer
3. Cumulative adjusted contract margin Answer
Answer #1

As per IFRS 15, para 35 "An entity transfers control of a good or service over time and, therefore, satisfies a performance obligation and recognises revenue over time, if one of the following criteria is met:

(c) the entity’s performance does not create an asset with an alternative use to the entity (see paragraph 36) and the entity has an enforceable right to payment for performance completed to date (see paragraph 37)."

In accordance with paragraph 37, an entity has a right to payment for performance completed to date if the entity would be entitled to an amount that at least compensates the entity for its performance completed to date in the event that the customer or another party terminates the contract for reasons other than the entity’s failure to perform as promised. An amount that would compensate an entity for performance completed to date would be an amount that approximates the selling price of the goods or services transferred to date (for example, recovery of the costs incurred by an entity in satisfying the performance obligation plus a reasonable profit margin) rather than compensation for only the entity’s potential loss of profit if the contract were to be terminated.

As per the question the equipment is highly specialized and has no alternative uses. Also any costs incurred by Supplier Corp. plus a specified profit margin will be paid to Supplier Corp. in the event of a contract cancellation hence Supplier Corp also has right to payment as per the above para of IFRS 15. Accordingly the condition of para 35(c) is met and hence revenue would be recognised over time. In the current case revenue would be recognised using input method.

a. Calculate (1) recognized revenue, (2) the gross profit, and (3) adjusted contract margin to be recorded.

2020 2021
(1) Recognized revenue for the year $814,286 $719,047
(2) Gross profit for the year $244,286 $139,047
(3) Adjusted contract margin for the year 20% 20%

Refer calculation below

b. Calculate (1) cumulative recognized revenue, (2) cumulative gross profit, and (3) cumulative adjusted contract margin at the end of 2021.

(1) Cumulative recognized revenue $1,533,333
(2) Cumulative gross profit $383,333
(3) Cumulative adjusted contract margin 20%

Refer calculation below

Particulars 2020 2021
Contract price $2,000,000 $2,000,000 (a)
Estimate of total costs at the end of the year $1,400,000 $1,500,000 (b)
Actual costs incurred till the end of the year $650,000
Actual costs incurred till the end of the year excluding inefficiencies $570,000 $1,150,000 (c)
Percentage of completion till the end of the year 40.71% 76.67% (d=c/b)
(1) Cumulative recognized revenue $814,286 $1,533,333 (e=a*d)
(2) Cumulative gross profit $244,286 $383,333 (f=e-c)
(3) Cumulative adjusted contract margin 20% 20% (g=(f-80000)/e)
Particulars 2020 2021 Remarks
(1) Recognized revenue for the year $814,286 $719,047 (1533333-814286=719,047)
(2) Gross profit for the year $244,286 $139,047 (383333-244286=139,047)
(3) Adjusted contract margin for the year 20% 19% (139047 / 719047=19%)

PS: If the answer is helpful a thumbs up would be appreciated